You’re a project engineer and want to figure out the internal rate of return on switching to a new process with new equipment. If this new equipment requires an initial investment of $5,500 but generates a net profit of $1,325 per year for the following 6 years, then what is the project's internal rate of return? Use interpolation methods or a spreadsheet to calculate your answer. Express your answer as a percentage value rounded to the nearest tenth of a percent (e.g. if you believe the answer to be 1.2%, enter 1.2). You may want to use linear interpolation.
Answer #1
IRR is that discount rate for which NPV = 0 or where
PV of cash inflows equals PV of cash outflows.
Hence,
5500 = 1325*PVIFA(IRR,6)
Solving for IRR
PVIFA(IRR,6) = 5500/1325 = 4.1509
The annuity interest factor for n = 6
and interest of 11% is 4.2305
The annuity interest factor for n = 6
and interest of 12% is 4.1114
Hence, IRR = 11%+1%*(4.2305-4.1509)/(4.2305-4.1114) = 11.67%